OpenStudy (anonymous):

I'm new to openstudy! Not really sure how this works but I seriously need some help with math! PLEASE!

TheSmartOne (thesmartone):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 @tiny.blob Posting your question can get you help faster :)

OpenStudy (is3535):



OpenStudy (anonymous):

When he was 35, Mr. Ramos bought 300 shares of stock for $56 per share in a well-established company. He received annual dividends in the amount of $1.25 per share. Upon retiring at age 65, he sold his shares at a price of $68 per share. What was Mr.Ramos' return on investment?

OpenStudy (anonymous):

Can you help me? @TheSmartOne

OpenStudy (is3535):

300 - 1.25 -56-65

OpenStudy (anonymous):

177.75 ? @is3535

TheSmartOne (thesmartone):

So he bought 300 stocks for $56 each. So \(\sf 300 \times 56=?\) That is his investment.

OpenStudy (anonymous):

16,800

OpenStudy (is3535):

so your answer is 177.75

TheSmartOne (thesmartone):

And per each stock he gets $1.25 per year And so 65 - 35 is 30 years so: So what is \(\sf 300 \times 1.25 \times 30 =?\)

TheSmartOne (thesmartone):

I doubt that is the answer, and it is that simple @is3535

OpenStudy (anonymous):

300*1.25*30= 11,250

TheSmartOne (thesmartone):

And then he sold his 300 shares each for $68 so what is \(\sf 300 \times 68=?\))

TheSmartOne (thesmartone):

Ignore that parenthesis, it's a typo :P

OpenStudy (is3535):

so im wrong

OpenStudy (anonymous):

20,400

TheSmartOne (thesmartone):

So his investment was \(\bf $16,800\) And he got back from this investment: \(\bf $11,250+ $20,400\) What is that equal to ^^

OpenStudy (anonymous):

$31,650?

TheSmartOne (thesmartone):

Correct. And now to find his return on invstement, we just subtract how much he made off of his investment by how much he invsted. So: \(\bf $31,650 - 16,800=?\)

OpenStudy (anonymous):

14,850

TheSmartOne (thesmartone):

And that is your final answer :) \(\Huge\bf\boxed{14,850}\)

OpenStudy (anonymous):

Thank you so much!

TheSmartOne (thesmartone):

Anytime :D :D :D

OpenStudy (anonymous):

do you have Mr. Ramos for your math teacher???

OpenStudy (anonymous):

No

OpenStudy (anonymous):

oh. well i have a teacher named mr ramos lol